OpenStudy (mathmath333):

Tom and wingspan are running the same race , the probablity of their winning are 1/5 and 1/2 respectively. Find the probablity that either of them will win the race.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ Tom and wingspan are running the same race , the probablity }\hspace{.33em}\\~\\ & \normalsize \text{ of their winning are 1/5 and 1/2 respectively. Find the }\hspace{.33em}\\~\\ & \normalsize \text{ probablity that either of them will win the race.}\hspace{.33em}\\~\\ & a.)\ \dfrac{7}{10} \hspace{.33em}\\~\\ & b.)\ \dfrac{3}{10} \hspace{.33em}\\~\\ & c.)\ \dfrac{1}{5} \hspace{.33em}\\~\\ & d.)\ \dfrac{7}{9} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (nuccioreggie):

CAN YOU UNBLOCK ME PLEASE

ganeshie8 (ganeshie8):

Since they are in same race, notice that, they cannot both win the race at the same time. So the events are mutually exclusive

OpenStudy (mathmate):

Events: T=tom wins W=wingspan wins Probability of either one of them winning is the probability of the union of the two events, \(T\cup W\), and the probability is P(\(T\cup W\))=P(T)+P(W)-P(\(T\cap W\)) if we assume events T and W are mutually exclusive (i.e. they cannot be cochampions), then P(\(T\cap W\))=0, and P(T∪W)=P(T)+P(W)-0 = ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & P(A\cup B)=P(A)+P(B) \hspace{.33em}\\~\\ \end{align}}\)